Listed below are foot lengths in inches for 11 randomly selected people taken in 1988. Find the​ range, variance, and standard deviation for the given sample data. Include appropriate units in the results. Are the statistics representative of the current population of all​ people?
8.6,9.2,8.8,9.3,8.5,9.9,10.2,8.910.2,10.5,9.7
The range of the sample data is
1
Expert's answer
2020-11-16T19:19:18-0500

Range = the difference between the highest and lowest numbers = 10.5 - 8.5 = 2 inches

Variance = ("\sum"(X2)-("\sum"X)2/N)/N = (984.42-103.82/11)/11 = 0.413 inches2

Standard deviation = (variance)0.5 = 0.643 inches

Too few people were chosen. Sample is not representative.